Twenty ounces of a 50% gold alloy are mixed with 80 oz of a 30% gold alloy. Find the percent concentration of the resulting gold alloy.
Find the cost per ounce of a gold alloy made from 25 oz of pure gold that costs $1286 per ounce and 45 oz of an alloy that costs $920 per ounce. (Round your answer to the nearest cent.)
Find the cost per pound of a "house blend" of coffee that is made from 16 lb of Central American coffee that costs $8 per pound and 34 lb of South American coffee that costs $5.50 per pound.
Find the cost per ounce of a mixture of 280 oz of a cologne that costs $6.50 per ounce and 420 oz of a cologne that costs $6.00 per ounce.

Answer by stanbon(75887)

Twenty ounces of a 50% gold alloy are mixed with 80 oz of a 30% gold alloy. Find the percent concentration of the resulting gold alloy.
-------
gold + gold = gold
0.50*20 + 0.30*80 = x(20+80)
-----
50*20 + 30*80 = 10000x
----
1000 + 2400 = 10000x
10000x = 3400
x = 0.34 or 34%
----------------
Cheers,
Stan H.
----------------

Find the cost per ounce of a gold alloy made from 25 oz of pure gold that costs $1286 per ounce and 45 oz of an alloy that costs $920 per ounce. (Round your answer to the nearest cent.)
Find the cost per pound of a "house blend" of coffee that is made from 16 lb of Central American coffee that costs $8 per pound and 34 lb of South American coffee that costs $5.50 per pound.
Find the cost per ounce of a mixture of 280 oz of a cologne that costs $6.50 per ounce and 420 oz of a cologne that costs $6.00 per ounce.